A Dialogue on the Best Form of Government for France

Duc Louis-Philippe d’Orleans ^{∗}

Feb 22, 1808

WMF: Good morning oh great and wise one.

Socrates: Good morning Weak-Minded Fool. And what have you come to ask me about?

WMF: I have been listening to the many philosophical debates Socrates, and I think I know which form of government is best for the great nation of France.

Socrates: And what philosophy, Mnsr. Fool, do you suggest make the best form of government for France?

WMF: Why, Socialism of course. For everybody in this system is equal, with benevolent leaders to ensure the equality of everyone.

Socrates: But Mnsr. Under this system can we ensure that the leaders are benevolent?

WMF: Of course, you are right socrates. A selfish or powerful outside force could easily step in and grab everything. There is no way to stop an unwilling participant from not sharing in everything he does. The followers of socialism must be truly good people at heart, but with beliefs of a benevolent society that cannot be ensured. But then, Democracy has to be the right form of government for France!

Socrates: Slow down Mnsr. Fool. Have we seen Democracy work?

WMF: Well Monsieur, we have not seen democracy work in the style that is being proposed today, but Athens was a good democracy.

Socrates: But who was Athens ruled by?

WMF: Athens was ruled by a smaller group of people who had citizenship status, not by everyone in the city. Not by women, not by the residents. And it did not have a population as large as France’s. While Democracy serves a small local government well, a democracy cannot work with as many citizens as France has. There will be too many people pulling in too many directions, there will be no clear representative of the people for foreign governments to negotiate with or to fight wars against. This system will either degenerate into total anarchy, or worse, a powerful but not necessarily benevolent leader may seize power fairly easily in this system. I can’t believe I fell for their irresponsible rhetoric. But I have been listening to the republicans, and they believe in creating a small group chosen by the people to represent them. Is this not a good thing for France?

Socrates: Perhaps it might be. Have we seen a system similar to a republican system in France?

WMF: Well, for several years now, we have had a directory which was a small group of representatives elected by the citizenry to rule them and pass laws for them.

Socrates: And is this a good system?

WMF: There is no cohesiveness. A group can easily seize power, and we don’t have balance.

∗With thanks to the great philosopher Plato

This is not a good system. Well, then, how about Enlightened Despotism? Surely a good strong leader is a great thing for France? Letting the people choose the best among themselves to govern them must be a good thing for France.

Socrates: I believe you are on the right track Monsieur, but let me ask you a question: Can the people err?

WMF: Of course the people can make mistakes. We are not perfect monsieur.

Socrates: And what happens if the people make a mistake monsieur?

WMF: Well, I suppose we have a bad form of government.

Socrates: Exactly, but then, if the people can err, can we trust them to make this choice?

WMF: Well, I guess not Monsieur.. But there must be a way to choose this powerful leader.

Socrates: Well, who can choose without making a mistake?

WMF: But of course! If we let a higher power choose, we cannot go wrong! Surely we must let our divine leader choose the man who will govern us, though we must also allow for a system that involves the people. Is this not a Constitutional Monarchy?

Socrates: I congratulate you monsieur, for you are truly an enlightened being. There is only one form of government that is good for France. It is a Constitutional Monarchy. Throughout France’s history, a strong leader has been good for France. Allowing for a council of advisors and giving this king veto power over their decision is the best for of Government that France can have.